OpenStudy (anonymous):

You made $3000 on a part-time job last year. If you expect to make the same amount this year, and the inflation rate is three percent, what is your purchasing power? Answer $3000 $3090 $2910 $3100

OpenStudy (ranga):

With inflation at 3% (or 0.03), the purchasing power will be 0.97 (100-0.03) of the income: $3,000 * .97 = $2,910.

OpenStudy (ranga):

I meant to say 1 - 0.03 = 0.97

OpenStudy (anonymous):

thank you can you help me another one pls?

OpenStudy (anonymous):

oh

OpenStudy (ranga):

I can try one more.

OpenStudy (anonymous):

Which of the following salaries will have a purchasing power of $24,035 if the inflation rate is five percent? Answer $25,815 $27,523 $26,000 $25,300

OpenStudy (ranga):

Inflation = 5% Purchasing power = 100 - 5 = 95% or 0.95 of the salary Let x be the salary. Purchasing power = 0.95x = 24,035 x = 24,035 / 0.95 = $25,300.

OpenStudy (anonymous):

oh okay i get it now thanks.

OpenStudy (ranga):

you are welcome.